# AI Foundation Models in Remote Sensing: A Survey

### v1

@@ V-B Domain-Specific Tasks
Domain-specific tasks in remote sensing refer to applications that address particular real-world challenges by leveraging the capabilities of foundation models. These tasks include environmental monitoring, archaeology, agriculture and forestry, urban planning and development, and disaster management. To highlight the versatility of foundation models in remote sensing, we present figure 5 that categorizes models based on their applicability to various domain-specific tasks, as well as computer vision tasks. This figure serves as a quick reference for researchers to identify suitable models for their specific needs.
@@
Figure 5: This diagram illustrates various domain-specific tasks in computer vision and their connections. Key areas include Environmental Monitoring, Agriculture and Forestry, Urban Planning and Development, Disaster Management, and Archaeology. Each domain comprises specific tasks such as change detection and object Detection. The relationships between these tasks and their application domains are depicted through connecting arrows, emphasizing the interconnected nature of computer vision applications across different fields.
  1. 1.
    Environmental Monitoring:
@@
    Environmental monitoring involves using remote sensing models to track changes in the environment, such as deforestation, desertification, and pollution. These models help in understanding the impact of human activities and natural phenomena on the environment.
  2. 2.
    Archaeology:
    In archaeology, remote sensing models have been used to identify and analyze archaeological features and sites. These models help in detecting features such as ruins, artifacts, and ancient structures from satellite imagery. CNNs and ViTs are often utilized for their ability to process high-resolution images and capture fine details.
  3. 3.
    Agriculture and Forestry:
@@
    In agriculture, remote sensing models are used to monitor crop health, estimate yields, and manage agricultural practices. These models help in optimizing resource use and improving agricultural productivity.
  4. 4.
    Urban Planning and Development:
@@
    In urban planning and development, remote sensing models are used to monitor and analyze urban expansion, infrastructure development, and land use changes. These models help in managing urban growth, planning new developments, and assessing the impact of urbanization.
  5. 5.
    Disaster Management:
@@
    Remote sensing models play a crucial role in disaster management by providing timely information on affected areas. These models are used to detect and assess damage from natural disasters such as earthquakes, floods, and hurricanes.
VI Discussion
@@ VI-A Synthesis of Findings
VI-A1 Model Performance.
  • Classification. The performance of foundation models on the BigEarthNet dataset [60] for classification tasks shows variations in accuracy, as presented in table VIII. Overall, msGFM [25] has the top performance of 92.90% (mAP), followed closely by SkySense [24] with a performance of 92.09%. Other notable performers include DeCUR [72], which achieved an mAP of 89.70%, and DINO-MC [77], with an mAP of 88.75%. SeCo [46] also demonstrated strong performance with an mAP of 87.81%, while DINO-MM [73] reached an mAP of 87.10%. On the other hand, models like CACo [47] and FoMo-Bench [5] have mAP of 74.98% and F1-Score of 68.33% respectively, showing competitiveness but room for improvement.
@@
  • Segmentation. For the segmentation tasks, we compared 12 foundation models which have been tested on the ISPRS Potsdam dataset. As shown in table VIII, SkySense [24] has the better performance out of all 12 models, with a mF1 Score of 93.99%. CMID [49] stands out with the highest mIoU of 87.04%, demonstrating its superior capability in accurately segmenting different regions within the dataset. For Overall Accuracy performance, BFM [7] has the highest OA score of 91.82%. Cross-Scale MAE [62], UPetu [18] and RSP [69] have mIoU scores of 76.17%, 83.17% and 65.30% respectively, showing competitive segmentation capabilities. GeoKR [39] reaches an mIoU of 70.48%, indicating robust segmentation performance but with room for improvement compared to CMID [49]. TOV scores the lowest mIoU at 60.34%, suggesting it may struggle with finer segmentation tasks compared to the other models.
@@
    BigEarthNet MMEarth [51] 78.6 mAP
    TABLE VII: This table provides an overview of the performance metrics for various models applied to the BigEarthNet dataset [60] for classification tasks. The performance is measured using mAP (Mean Average Precision) and F1 Score.
@@
    ISPRS Potsdam BFM [7] 92.58 OA
    TABLE VIII: This table provides an overview of the performance metrics for various models applied to the ISPRS Potsdam[30] dataset for segmentation tasks. The performance is measured using Mean Intersection over Union (mIoU) and Overall Accuracy (OA).
  • Object Detection. In table X, the performance of foundation models on the DOTA, DIOR, and DIOR-R datasets are evaluated based on their Mean Average Precision (mAP) and Average Precision at 50% (AP50).
@@
  • Change Detection. For the change detection tasks, we compared the performance of foundation models on the OSCD and LEVIR-CD datasets. The models were evaluated based on their F1 Scores, which provide a balanced measure of precision and recall. As shown in the table, the performance varies significantly across different models and datasets.
    In contrast, the LEVIR-CD dataset reveals higher performance metrics across the models. MTP [70] achieves the highest F1 Score of 92.67% and SkySense [24] follows closely with an F1 Score of 92.58%, demonstrating their robust performance. SWiMDiff reaches a lower F1 Score of 80.90% compared to its peers but still indicates effective performance in the LEVIR-CD dataset.
@@
    TABLE IX: This table provides an overview of the performance metrics for various models applied to the DOTA[16, 79, 15] dataset, DIOR[38] and DIOR-R[9] dataset for object detection task. The performance is mainly measured using Mean Average Precision (mAP)
    Dataset Model Performance (%) Metrics
    SeCo [46] 46.94 F1 Score
    MATTER [1] 49.48 F1 Score
    CACo [47] 52.11 F1 Score
    GFM [48] 59.82 F1 Score
    SWiMDiff [65] 49.60 F1 Score
    SpectralGPT [28] 54.29 F1 Score
    SkySense [24] 60.06 F1 Score
    DINO-MC [77] 52.71 F1 Score
    OSCD MTP [70] 53.36 F1 Score
    RSP [69] 90.93 F1 Score
    RingMo [61] 91.86 F1 Score
    RIngMo-lite [76] 91.56 F1 Score
    SwiMDiff [65] 80.90 F1 Score
    SkySense [24] 92.58 F1 Score
    UPetu [18] 88.50 F1 Score
    LEVIR-CD MTP [70] 92.67 F1 Score
@@
    TABLE X: This table provides an overview of the performance metrics for various models applied to the Onera Satellite Change Detection (OSCD) dataset[6] and the LEVIR-CD dataset[8] for change detection downstream tasks. The performance is measured using F1-Score.
@@
  • Influence of Pre-training Methods. Various pre-training methods have a substantial impact on the performance of foundation models in remote sensing. Models pre-trained using self-supervised learning (SSL) techniques, such as contrastive learning (CL) and masked autoencoders (MAE), consistently exhibit superior performance compared to those pre-trained with traditional supervised learning. For instance, models like SkySense [24] and SeCo [46], which employ contrastive learning, achieve higher accuracy and robustness in tasks like scene classification and object detection. Additionally, models utilizing predictive coding techniques, such as SatMAE [11] and Scale-MAE [55], demonstrate enhanced capability in handling multi-temporal and multi-spectral data, leading to significant improvements in change detection and environmental monitoring applications. These findings highlight the critical role of innovative pre-training methods in maximizing the effectiveness of foundation models and suggest that continued exploration and refinement of these techniques are essential for advancing remote sensing capabilities.
VI-A2 Practical Implications
The advancements in foundation models have significant practical implications across various domains:

### v2

@@ VI-B Tasks
Different applications in remote sensing address particular real-world challenges by leveraging the capabilities of foundation models. These tasks include environmental monitoring, archaeology, agriculture, urban planning and development, and disaster management. To highlight the versatility of foundation models in remote sensing, we present table III that categorizes models based on their applicability to various applications, as well as the different image analysis methods used. This figure serves as a quick reference for researchers to identify suitable models for their specific needs.
Image Analysis by Levels
Tasks Image-Level Pixel-Level Region-Level Spatial-Temporal Related Work
Environmental Monitoring Land Cover Change Detection
$\checkmark$
[66, 2, 96, 87, 70, 67, 69, 109, 91, 40]
[32, 24, 111, 99, 106, 49, 59, 122, 95]
Deforestation Monitoring
$\checkmark$
[56, 2, 6, 96, 79, 87, 57, 45, 100, 16, 89]
[70, 67, 69, 7, 119, 78, 109, 102, 27]
[108, 28, 88, 26, 42, 22, 40, 92, 32, 24]
[33, 118, 99, 123, 106, 49, 110, 122, 95, 117]
Water Body Analysis
$\checkmark$
$\checkmark$
$\checkmark$
[96, 87, 70, 27, 32, 99, 49, 95]
Forest Cover mapping
$\checkmark$
$\checkmark$
[96, 87, 67, 69, 109, 27, 40, 106, 49]
[59, 122, 95]
Biomass Estimation
[77]
Weather/Climate Prediction
$\checkmark$
[82, 119, 101]
Cloud Removal [33]
Moisture Content Measurement [92]
Agriculture Crop Type Mapping
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
[96, 87, 70, 27, 32, 99, 49, 95, 83]
Weed Detection
$\checkmark$
[56, 6, 96, 87, 57, 124, 100, 89, 70, 27]
[8, 32, 99, 123, 49, 110, 95, 22]
Disease Monitoring
$\checkmark$
$\checkmark$
$\checkmark$
[56, 6, 96, 87, 57, 100, 89, 70, 27]
[32, 99, 123, 49, 110, 95, 22]
Forecasting [82, 119]
Soil Parameter Estimation [117]
Yield Estimation
$\checkmark$
$\checkmark$
[56, 2, 6, 96, 79, 87, 57, 45, 100, 16, 89]
[70, 67, 69, 7, 78, 109, 102, 27]
[108, 88, 26, 42, 40, 32, 24, 33, 118]
[99, 123, 106, 122, 95, 117]
Agricultural Pattern Segmentation
$\checkmark$
[66]
Archaeology
Artifact Classification
and Recognition
$\checkmark$
$\checkmark$
[84, 66, 56, 2, 6, 96, 105, 79, 87]
[124, 100, 16, 89, 70, 67, 77, 69, 7, 78, 109]
[27, 108, 44, 88, 26, 82, 42, 91, 58]
[8, 32, 24, 33, 111, 118, 99, 72, 123, 106]
[45, 57, 117, 95, 102, 40, 59, 122]
Detection of Archaeological Structures
$\checkmark$
[56, 6, 96, 87, 57, 124, 100, 89, 70, 27]
[8, 32, 99, 123, 49, 110, 95, 22]
Semantic Segmentation
$\checkmark$
[56, 2, 6, 96, 79, 87, 57, 45, 100, 16, 89]
[70, 67, 69, 7, 119, 78, 109, 102, 27]
[108, 28, 88, 26, 42, 22, 40, 92, 32, 24]
[33, 118, 99, 123, 106, 49, 110, 122, 95, 117]
Texture/Structural Analysis [2]
Pattern Recognition
$\checkmark$
$\checkmark$
[56, 6, 96, 87, 57, 100, 89, 70, 27]
[32, 99, 123, 49, 110, 95, 22]
Urban Planning & Development Traffic Monitoring
$\checkmark$
$\checkmark$
$\checkmark$
[96, 87, 70, 27, 32, 99, 49, 95]
Land Cover/Use Classification
$\checkmark$
[84, 66, 56, 2, 6, 96, 105, 79, 87]
[124, 100, 16, 89, 70, 67, 77, 69, 7, 78, 109]
[27, 108, 44, 88, 26, 82, 42, 91, 58]
[8, 32, 24, 33, 111, 118, 99, 72, 123, 106]
[45, 57, 117, 95, 102, 40, 59, 122, 83]
Road Crack Detection
$\checkmark$
[56, 6, 96, 87, 57, 124, 100, 89, 70, 27]
[8, 32, 99, 123, 49, 110, 95, 22]
Air Quality Monitoring
$\checkmark$
$\checkmark$
[33, 56, 6, 96, 87, 57, 100, 89, 70, 27]
[32, 99, 123, 49, 110, 95, 22]
Building Extraction
$\checkmark$
$\checkmark$
[27]
Object/Video Tracking 1 c|
$\checkmark$
$\checkmark$
[119]
Infrastructure Monitoring [119, 44]
Disaster Management Landslide Risk Monitoring
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
[96, 87, 70, 27, 32, 99, 49, 95]
Disaster Response [117]
Real-Time Detection and Mapping
$\checkmark$
$\checkmark$
$\checkmark$
[96, 87, 70, 27, 32, 99, 49, 95]
Building Damage Assessment
$\checkmark$
$\checkmark$
$\checkmark$ [96, 87, 70, 27, 32, 99, 49, 95]
Critical Infrastructure Detection
$\checkmark$
$\checkmark$
[56, 6, 96, 87, 57, 124, 100, 89, 70, 27]
[8, 32, 99, 123, 49, 110, 95, 22]
Flood/Fire Mapping and Prediction
$\checkmark$
$\checkmark$
$\checkmark$
[96, 87, 67, 69, 109, 27, 40]
[106, 49, 122, 95]
Crowd and Vehicle Detection
$\checkmark$
$\checkmark$
[66, 2, 96, 87, 70, 67, 69, 109, 91, 40]
[32, 24, 111, 99, 106, 49, 59, 122, 95]
@@
TABLE III: This diagram illustrates various tasks in different applications for remote sensing. Key areas include Environmental Monitoring, Agriculture, Urban Planning and Development, Disaster Management, and Archaeology. Each domain comprises specific tasks in different image analysis levels like image-level, pixel-level, region-level, and spatial-temporal. The relationships between these tasks and their applications are depicted through checkmarks, emphasizing the interconnected nature of image analysis methods across different fields.
VI-B1 Environmental Monitoring
@@
According to Himeur et al. (2022), environmental monitoring utilizes remote sensing models to observe and track environmental changes, including deforestation, desertification, and pollution. These models play a crucial role in analyzing the effects of human activities and natural phenomena on the environment [39].
VI-B2 Agriculture
@@
In agriculture, remote sensing models are used to monitor crop health, estimate yields, and manage agricultural practices. According to Kamilaris et al. (2018), these models help optimize resource use and improve agricultural productivity [52].
VI-B3 Archaeology
In archaeology, remote sensing models have been used to identify and analyze archaeological features and sites. According to Argyrou et al. (2022), these models help detect features such as ruins, artifacts, and ancient structures from satellite imagery, leveraging technologies like Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) to process high-resolution images and capture fine details [5]. Mantovan et al. (2020) also highlight the effectiveness of AI models, particularly CNNs, in locating challenging terrestrial archaeological sites and processing multispectral data [68].
VI-B4 Urban Planning and Development
@@
In urban planning and development, remote sensing models are used to monitor and analyze urban expansion, infrastructure development, and land use changes. According to Jha et al. (2021), these models play a critical role in managing urban growth, planning new developments, and assessing the impact of urbanization by providing essential data for smart city planning and sustainable development [48].
VI-B5 Disaster Management
@@
Remote sensing models play a crucial role in disaster management by providing timely information on affected areas. According to Abid et al. (2021), these models are used to detect and assess damage from natural disasters like earthquakes, hurricanes, and floods, enabling rapid response and recovery efforts [1].
VII Discussion
@@ VII-A Synthesis of Findings
VII-A1 Model Performance
In this section, we present the performance metrics of recent foundation models in remote sensing based on results reported in the original papers. All performance numbers mentioned here are sourced directly from the original studies to ensure accuracy and consistency in evaluating these models. These metrics provide insights into the models’ effectiveness across tasks like semantic segmentation, object detection, and change detection, highlighting their strengths and limitations under different experimental setups.
  • Image-Level The performance of foundation models on the BigEarthNet dataset [85] for classification tasks shows variations in accuracy, as presented in table V. Overall, msGFM [33] has the top performance of 92.90% (mAP), followed closely by SkySense [32] with a performance of 92.09%. Other notable performers include DeCUR [102], which achieved an mAP of 89.70%, and DINO-MC [111], with an mAP of 88.75%. SeCo [66] also demonstrated strong performance with an mAP of 87.81%, while DINO-MM [105] reached an mAP of 87.10%. On the other hand, models like CACo [67] and FoMo-Bench [8] have mAP of 74.98% and F1-Score of 68.33% respectively, showing competitiveness but room for improvement.
@@ VII-A1 Model Performance
    The classification advancements observed in remote sensing models stem from sophisticated pre-training techniques that capture both spatial and spectral complexity across vast datasets. SkySense, for example, shows an average improvement of 2.76% over recent models by implementing multi-granularity contrastive learning on a diverse dataset of 21.5 million optical and SAR sequences [32]. This approach enables SkySense to learn nuanced spatial and temporal relationships across modalities, enhancing generalization in varied environmental conditions. Such multi-granular representation proves crucial in remote sensing, where scene classification often depends on subtle spectral differences that simpler models may overlook. Likewise, HyperSIGMA [95], pre-trained on the expansive HyperGlobal-450K hyperspectral dataset [95], leverages its sparse sampling attention mechanism to optimize spectral-spatial feature extraction in high-dimensional hyperspectral data. By selectively focusing on critical spectral bands and reducing redundancy, HyperSIGMA achieves high classification accuracy across hyperspectral scenes, a marked improvement over previous models that struggled with hyperspectral data complexity.
    These models highlight the importance of designing pre-training strategies that capture multi-modal features and effectively utilize dataset diversity, as these elements directly impact the robustness and accuracy of classification in remote sensing applications.
  • Pixel-Level For the segmentation tasks, we compared 12 foundation models which have been tested on the ISPRS Potsdam dataset. As shown in table V, SkySense [32] has the better performance out of all 12 models, with a mF1 Score of 93.99%. CMID [70] stands out with the highest mIoU of 87.04%, demonstrating its superior capability in accurately segmenting different regions within the dataset. For Overall Accuracy performance, BFM [11] has the highest OA score of 91.82%. Cross-Scale MAE [88], UPetu [24] and RSP [96] have mIoU scores of 76.17%, 83.17% and 65.30% respectively, showing competitive segmentation capabilities. GeoKR [56] reaches an mIoU of 70.48%, indicating robust segmentation performance but with room for improvement compared to CMID [70]. TOV scores the lowest mIoU at 60.34%, suggesting it may struggle with finer segmentation tasks compared to the other models.
@@
    MMEarth [72] 78.6 mAP
    BigEarthNet [85] Shallow CNN* [85] 70.98 F1-Score
    TABLE IV: This table provides an overview of the performance metrics for various models applied to the BigEarthNet dataset [85] for image-level tasks. The performance is measured using mAP (Mean Average Precision) and F1 Score. *Performance for the shallow CNN model are sourced from the original BigEarthNet [85] paper.
@@
    BFM [11] 92.58 OA
    ISPRS Potsdam R-SegNet* [127] 91.37 OA
    TABLE V: This table provides an overview of the performance metrics for various models applied to the ISPRS Potsdam[43] dataset for pixel-level tasks. The performance is measured using Mean Intersection over Union (mIoU) and Overall Accuracy (OA). *Non-FM
    For the change detection tasks, we compared the performance of foundation models on the OSCD and LEVIR-CD datasets. The models were evaluated based on their F1 Scores, which provide a balanced measure of precision and recall. As shown in the table, the performance varies significantly across different models and datasets.
    In contrast, the LEVIR-CD dataset reveals higher performance metrics across the models. MTP [99] achieves the highest F1 Score of 92.67% and SkySense [32] follows closely with an F1 Score of 92.58%, demonstrating their robust performance. SWiMDiff reaches a lower F1 Score of 80.90% compared to its peers but still indicates effective performance in the LEVIR-CD [12] dataset.
  • Region-Level In table VII, the performance of foundation models on the DOTA, DIOR, and DIOR-R datasets for object detection are evaluated based on their Mean Average Precision (mAP) and Average Precision at 50% (AP50).
@@
    TABLE VI: This table provides an overview of the performance metrics for various models applied to the DOTA[21, 113, 20] dataset, DIOR[55] and DIOR-R[14] dataset for region-level task. The performance is mainly measured using Mean Average Precision (mAP). *Model performance are aquired from original dataset papers. AOPG is Anchor-free Oriented Proposal Generator.
    Dataset Model F1 Score
    SeCo [66] 46.94
    MATTER [2] 49.48
    CACo [67] 52.11
    GFM [69] 59.82
    SWiMDiff [91] 49.60
    SpectralGPT [40] 54.29
    SkySense [32] 60.06
    DINO-MC [111] 52.71
    HyperSIGMA [95] 59.28
    MTP [99] 53.36
    OSCD [10] CNNs* [10] 89.66 (OA)
    RSP [96] 90.93
    RingMo [87] 91.86
    RIngMo-lite [109] 91.56
    SwiMDiff [91] 80.90
    SkySense [32] 92.58
    UPetu [24] 88.50
    LEVIR-CD[12] STANet* [12] 85.4
@@
    TABLE VII: This table provides an overview of the F1 Score for various models applied to the Onera Satellite Change Detection (OSCD) dataset[10] and the LEVIR-CD dataset[12] for spacial-temporal downstream tasks. *Performance for the models are sourced from the original dataset papers. STANet is Spatial-Temporal Attention Network.
@@
  • Influence of Pre-training Methods. Various pre-training methods have a substantial impact on the performance of foundation models in remote sensing. Models pre-trained using SSL techniques, such as contrastive learning (CL) and masked autoencoders (MAE), consistently exhibit superior performance compared to those pre-trained with traditional supervised learning. For instance, SkySense which uses a multi-granularity contrastive learning approach, outperforms other models by approximately 3.6% in scene classification and object detection tasks [32]. Similarly, Seco, based on seasonal contrast learning, yields superior performance for land-cover classification, improving metrics by up to 7% over ImageNet-pre-trained models[66] . In handling multi-temporal and multi-spectral data, models like SatMAE [16] and Scale-MAE [78], using masked autoencoding, achieve improvements in change detection, with SatMAE showing up to a 14% performance gain in land cover classification [16] and Scale-MAE offering a 1.7% mIoU improvement for segmentation across varied resolutions [78]. These findings highlight the critical role of innovative pre-training methods in maximizing the effectiveness of foundation models and suggest that continued exploration and refinement of these techniques are essential for advancing remote sensing capabilities.
    Foundation models like SatMAE, RingMo, A2-MAE, and ORBIT each demonstrate strong performance, but practical trade-offs are essential to consider, especially for application-specific constraints [16, 87, 122, 101]. SatMAE, based on a transformer architecture, effectively leverages temporal and multi-spectral embeddings to capture complex spatiotemporal patterns in satellite imagery. This strength, however, comes at the cost of significant computational requirements, which may not be feasible for real-time monitoring applications in resource-constrained environments.
    In contrast, RingMo provides a more lightweight vision transformer architecture, offering efficient model inference and a balance between performance and computational demands. This makes RingMo particularly suitable for rapid-inference tasks like disaster response monitoring, where real-time processing is critical. [87] A2-MAE introduces an anchor-aware masking strategy, optimizing spatial-temporal-spectral representations and allowing effective integration of multi-source data. This design enhances its adaptability to varied data resolutions and modalities, yet the model’s complex encoding techniques add to its computational load, suggesting a fit for applications that require high accuracy over efficiency [122].
    Finally, ORBIT, designed with 113 billion parameters, is exceptionally scalable, achieving high-throughput performance for Earth system predictability tasks. While it excels in large-scale predictive tasks, the model’s considerable resource requirements limit its deployment to specialized high-performance computing environments. [101] These trade-offs highlight the importance of selecting a model that aligns with specific operational goals, whether for maximizing accuracy or minimizing computational overhead.
    Furthermore, recent studies comparing SSL approaches highlight the distinct advantages of generative methods like Masked Autoencoders (MAE) over contrastive methods for time-series data, especially when labeled data is limited[61]. Unlike contrastive approaches that emphasize distinguishing between similar and dissimilar pairs, generative methods such as MAE reconstruct data from masked segments, allowing them to capture complex underlying structures and relationships within the data. This reconstruction-based learning proves particularly advantageous for time-series and multi-spectral applications in remote sensing, where temporal and spectral dependencies are essential. Consequently, MAE-based models can achieve stronger representations under sparse labeling conditions, positioning them as powerful tools for remote sensing tasks that require nuanced temporal analysis.
VII-A2 Practical Implications
Foundation models offer transformative capabilities in remote sensing by building upon established applications like multi-spectral and time-series data analysis. While these applications have traditionally relied on machine learning and deep learning, foundation models reduce the need for labeled data and enable rapid adaptation to new tasks, providing robust solutions in areas previously limited by data constraints and task-specific architectures. Consequently, the advancements in foundation models have significant practical implications across various areas: